# SketchGuard: Scaling Byzantine-Robust Decentralized Federated Learning via Sketch-Based Screening

### v2

@@ 5.2.1. Attack Strategy Resilience
Figure 3. Impact of fraction of malicious clients on TER across different datasets and attack types.
Figure 3 evaluates SketchGuard’s robustness against both directed deviation and Gaussian attacks on FEMNIST and CelebA datasets, averaged across all network topologies. SketchGuard achieves robustness equivalent to state-of-the-art methods BALANCE and UBAR across both attack strategies. Compared to BALANCE, SketchGuard exhibits statistically identical performance with mean absolute TER deviation of 0.02 percentage points and maximum deviation of 0.49 percentage points. Compared to UBAR, SketchGuard shows mean absolute TER deviation of 0.35 percentage points and maximum deviation of 1.51 percentage points, with SketchGuard having marginally higher TER on average by 0.20 percentage points; however, UBAR’s loss-based screening mechanism requires significantly higher computational overhead than both BALANCE and SketchGuard, as demonstrated in Section 5.3. Against both directed deviation (Figures 3a, 3c) and Gaussian attacks (Figures 3b, 3d), SketchGuard maintains consistently low error rates across all Byzantine client fractions, demonstrating that sketch-based compression preserves the robustness properties of full-precision similarity-based defenses while enabling substantial efficiency gains.
5.2.2. Sketch Size Robustness
SketchGuard demonstrates remarkable robustness to sketch size selection. Across $k$ values ranging from 500 to 100,000 on both FEMNIST and CelebA datasets under 50% Byzantine clients with directed deviation attacks, performance remains completely stable (TER = 15.59% for FEMNIST, 10.35% for CelebA across all $k$ values). Even with compression ratios exceeding 13,000:1 ($k=500$ for FEMNIST’s 6.6M parameters), SketchGuard preserves full Byzantine resilience. This insensitivity to $k$ stems from Count Sketch’s distance preservation guarantees: while smaller $k$ values introduce larger approximation errors, the resulting sketch-based distances remain sufficiently accurate to distinguish Byzantine node signatures from honest node behavior. The adaptive threshold mechanism $\gamma_{\text{eff}}$ compensates for increased approximation error, ensuring robust filtering across the entire practical range of sketch sizes.
5.2.3. Topology-Dependent Robustness
Figure 4. Impact of fraction of malicious clients on TER across different network topologies.
Figure 4 presents SketchGuard’s robustness across five network topologies, averaged over attack types and datasets. SketchGuard demonstrates identical robustness to BALANCE and UBAR across all network topologies, from minimal connectivity in Ring topology to maximum connectivity in Fully Connected networks. In well-connected topologies (ER graphs with $p\geq 0.45$, Fully Connected), all three Byzantine-robust methods maintain low error rates below 20% even when 80% of the network is compromised, significantly outperforming KRUM and the non-robust D-FedAvg baseline. In extremely sparse topologies (Ring), BALANCE, UBAR, and SketchGuard exhibit graceful degradation, tolerating up to 40% Byzantine clients before substantial performance loss. Despite having only two neighbors per node in the Ring topology, SketchGuard maintains performance parity with full-precision methods across all attack intensities, demonstrating that sketch-based compression does not compromise filtering effectiveness even under severe connectivity constraints.
5.3. Computational Efficiency
Figure 5 evaluates SketchGuard’s computational efficiency through network size and model dimensionality scaling experiments on FEMNIST with 50% Byzantine clients under directed deviation attacks. Network scaling uses the baseline 6.6M parameter model across varying k-regular graph sizes, while model dimensionality scaling uses a fixed k-regular graph with node degree 154 across varying model sizes as detailed in Appendix C.3.
@@ 5.3. Computational Efficiency
Figure 5. Impact of network size and model dimensionality on per-node computation time.
5.3.1. Network Size Scalability
SketchGuard exhibits near-constant computational cost as network connectivity increases from 16 to 299 neighbors (Figure 5a), averaging 0.35s per round, while BALANCE and UBAR scale linearly with neighbor count. At smaller network sizes ($\leq$32 neighbors), SketchGuard’s sketch generation overhead dominates, resulting in higher absolute costs than full-precision methods. However, as connectivity increases beyond 96 neighbors, SketchGuard’s compressed screening operations yield substantial benefits. The efficiency gap widens multiplicatively with network size: at 154 neighbors, SketchGuard requires 0.34s compared to BALANCE’s 0.51s and UBAR’s 1.10s, achieving 33% and 69% savings respectively; at 299 neighbors, these savings expand to 60% and 82%, with SketchGuard requiring only 0.39s versus BALANCE’s 0.99s and UBAR’s 2.14s. This scalability advantage stems from neighbor screening operating in compressed $k$-dimensional space rather than full $d$-dimensional space, yielding multiplicative gains proportional to network connectivity.
5.3.2. Model Dimensionality Scalability
@@
SketchGuard demonstrates sub-linear scaling as model size increases from 220K to 60M parameters (Figure 5b), contrasting sharply with the linear scaling of BALANCE and UBAR. For models exceeding 26M parameters, SketchGuard maintains average computational improvement of 64% over state-of-the-art methods. At 60M parameters, SketchGuard requires 2.0s versus BALANCE’s 4.2s and UBAR’s 10.0s, achieving 52% and 80% reductions respectively. This sub-linear scaling emerges from Count Sketch’s property that sketch size $k$ depends on approximation parameters ($\epsilon$, $\delta$) rather than model dimension $d$. While full-precision methods incur $O(d|\mathcal{N}_{i}|)$ comparison costs that grow linearly with model size, SketchGuard’s sketch comparison cost remains fixed at $O(k|\mathcal{N}_{i}|)$ regardless of $d$, proving particularly valuable for large-scale DFL scenarios involving billion-parameter models (Borzunov et al., 2023).
5.4. Communication Efficiency
SketchGuard reduces per-round communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ denotes accepted neighbors post-filtering.
5.4.1. Compression Ratios and Bandwidth Reduction
With compression ratios exceeding 6300:1 for both datasets, sketch transmission overhead becomes negligible compared to full model exchanges. Table 3 presents communication efficiency under different filtering scenarios. In scenarios where Byzantine filtering successfully rejects 70% of malicious neighbors, SketchGuard achieves approximately 70% communication reduction relative to full-precision methods. Even in benign settings where all neighbors are accepted, the sketch overhead of $O(k|\mathcal{N}_{i}|)$ represents less than 0.02% additional cost compared to the $O(d|\mathcal{N}_{i}|)$ baseline due to the compression ratio $k/d\approx 1/6300$, introducing negligible bandwidth penalty.
Table 3. Communication Overhead Comparison
Scenario Accepted Neighbors Reduction
Benign (no filtering) $|\mathcal{S}_{i}|=|\mathcal{N}_{i}|$ <<0.02% penalty
50% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.5|\mathcal{N}_{i}|$ $\sim$50%
70% Byzantine filtering $|\mathcal{S}_{i}|\approx 0.3|\mathcal{N}_{i}|$ $\sim$70%
5.4.2. Multiplicative Scaling Benefits
The communication savings scale multiplicatively with both model dimensionality and network connectivity. Consider a fully connected network with 100 neighbors per node and FEMNIST’s 6.6M parameter model under 50% Byzantine clients, where Count Sketch compresses models to 1K parameters:
  • SketchGuard: Transmits 100K sketch parameters plus 330M full model parameters, totaling $\sim$330M parameters.
  • Full-Precision Methods: Transmit 660M parameters to all neighbors.
  • Reduction: $\sim$50% bandwidth savings.
These bandwidth savings prove critical for resource-constrained devices and bandwidth-limited networks characteristic of real-world decentralized learning deployments. As model sizes grow and networks expand, SketchGuard’s communication efficiency becomes essential for practical deployment.
6. Conclusions and Future Work
@@
We proposed SketchGuard, a framework for scaling Byzantine-robust decentralized federated learning through sketch-based neighbor screening. By decoupling Byzantine filtering from model aggregation, our approach reduces communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$ while preserving robustness guarantees of full-precision similarity-based defenses.
@@
Our theoretical analysis establishes that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds, where approximation errors introduce only a $(1+O(\epsilon))$ factor in the effective threshold parameter. We prove convergence guarantees in both strongly convex and non-convex settings, achieving optimal rates for their respective problem classes. Comprehensive experiments demonstrate that SketchGuard maintains identical robustness to state-of-the-art methods, with mean TER deviation of only up to 0.35 percentage points. This robustness equivalence is achieved while reducing computation time by up to 82%, with efficiency gains scaling multiplicatively with model dimensionality and network connectivity. Communication overhead scales proportionally with the fraction of accepted neighbors, achieving reductions of 50-70% under typical Byzantine attack scenarios.
SketchGuard addresses the fundamental scalability bottleneck limiting real-world deployment of robust decentralized learning. Future work includes efficient hash synchronization protocols for dynamic networks, adaptive sketch sizing based on attack intensity, and theoretical extensions to non-IID data distributions. As FL expands into web-scale deployments with untrusted participants, scalable robustness mechanisms like SketchGuard become essential infrastructure for maintaining both security and performance in distributed machine learning systems.
Acknowledgements.
This work is supported by the Australian Research Council (ARC) through Discovery Project grant DP240102088 and Linkage Infrastructure, Equipment and Facilities grant LE200100049.
Appendix A Proof of Theorem 1
@@ A.1. Setup and Notation
(10) $\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t}),$
@@
A.2. Model Update Analysis
@@
(11) $\quad+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t}$
@@
(12) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2}+\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
@@
(13) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$
@@
(14) $\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})$
A.3. Applying Smoothness
@@
(15) $\quad+\frac{L}{2}\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|^{2}$
@@
(16) $\quad\left.+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
A.4. Bounding the Quadratic Term
@@
(17) $\quad+2\left\|\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$
A.5. Impact of Sketch-Based Filtering
@@
(18) $\quad\leq\gamma\exp(-\kappa t/T)\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})\|$
@@
(19) $\quad\leq\gamma\sqrt{\frac{1+\epsilon}{1-\epsilon}}\exp(-\kappa t/T)\|\mathbf{w}_{i}^{t+1/2}\|$
@@
(20) $\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$
@@ A.6. Taking Expectation
(21) $\quad+L(1-\alpha)^{2}\gamma_{\text{eff}}^{2}\psi^{2}$
A.7. Simplifying with Parameter Constraints
@@
(22) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
@@
A.8. Applying Strong Convexity
@@
(23) $\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}\geq 2\mu(F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*}))$
@@
(24) $\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$
A.9. Telescoping and Final Bound
@@
(25) $\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$
This completes the proof of Theorem 1. The key observation is that the compression error only affects the convergence through $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, introducing a controllable degradation factor while maintaining the same convergence rate as the state-of-the-art.
Appendix B Proof of Theorem 2
@@
B.1. Starting from Smoothness

### v3

@@ 5.2 Byzantine Robustness Evaluation
Figure 2: Impact of fraction of malicious clients on TER across different datasets and attack types. SketchGuard maintains robustness equivalent to full-precision methods BALANCE and UBAR despite operating on compressed representations.
Figure 2 evaluates whether sketch-based compression degrades Byzantine robustness compared to full-precision defenses. Across all four attack types on FEMNIST, CelebA, and Sent140 datasets, SketchGuard maintains robustness statistically indistinguishable from BALANCE and UBAR, with mean absolute TER deviation of only 0.38 percentage points versus BALANCE and 0.50 percentage points versus UBAR. This confirms that Count Sketch’s distance preservation properties (Lemma 2.1) translate to equivalent filtering decisions in practice.
Against directed deviation and Krum attacks, all three similarity-based methods (BALANCE, UBAR, SketchGuard) maintain TER below 20% on FEMNIST and below 13% on CelebA even at 80% Byzantine clients, while D-FedAvg and KRUM collapse to near-random accuracy (TER >> 60%). The Krum defense fails against its targeted attack (TER >> 60% at 30%+ Byzantine), whereas similarity-based filtering—whether on full models or sketches—remains effective. For Gaussian attacks, SketchGuard matches BALANCE within 0.03 percentage points on average.
For backdoor attacks, SketchGuard achieves ASR comparable to BALANCE and UBAR across all datasets. On FEMNIST, all three methods achieve low ASR (6.98–9.94%), demonstrating that similarity-based filtering effectively identifies the scaled model updates characteristic of backdoor attacks. On CelebA and Sent140, all robust methods substantially outperform D-FedAvg. These results confirm that sketch compression does not create vulnerabilities to backdoor attacks.
5.3 Computational Efficiency
@@
Figure 3: Impact of network size and model dimensionality on per-node computation time. SketchGuard achieves substantial savings while maintaining equivalent robustness (Section 5.2).
Having established that SketchGuard matches state-of-the-art robustness, we now demonstrate the efficiency gains that motivate sketch-based filtering. Figure 3 evaluates computational scaling on FEMNIST with 50% Byzantine clients under directed deviation attacks.
SketchGuard exhibits near-constant cost as network connectivity increases from 16 to 299 neighbors (Figure 3a), averaging 0.35s per round, while BALANCE and UBAR scale linearly. At 299 neighbors, SketchGuard achieves 60% savings versus BALANCE (0.39s vs 0.99s) and 82% versus UBAR (0.39s vs 2.14s). This advantage grows with network density because sketch comparisons cost $O(k)$ per neighbor versus $O(d)$ for full-precision methods.
@@
For model dimensionality scaling (Figure 3b), SketchGuard demonstrates sub-linear growth from 220K to 60M parameters, while BALANCE and UBAR scale linearly. At 60M parameters, SketchGuard requires 2.0s versus 4.2s (BALANCE) and 10.0s (UBAR)—52% and 80% reductions. This sub-linear scaling emerges because sketch size $k$ depends on approximation parameters ($\epsilon$, $\delta$) rather than model dimension $d$, making SketchGuard increasingly valuable as models grow toward billion-parameter scales (Borzunov et al., 2023).
5.4 Communication Efficiency
SketchGuard reduces per-round communication from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$, where $|\mathcal{S}_{i}|\leq|\mathcal{N}_{i}|$ is the accepted neighbor count. With compression ratios of approximately 6600:1, sketch overhead is negligible (<<0.02% in benign settings). When Byzantine filtering rejects 50–70% of neighbors, communication reductions of 50–70% follow directly. These savings scale multiplicatively with model size and network connectivity, addressing the fundamental scalability bottleneck that has limited practical deployment of Byzantine-robust DFL.
5.5 Sensitivity Analysis
5.5.1 Sketch Size Robustness
A practical concern for sketch-based methods is sensitivity to the sketch size parameter $k$. SketchGuard demonstrates remarkable insensitivity: across $k$ values ranging from 500 to 100,000 on both FEMNIST and CelebA under 50% Byzantine clients with directed deviation attacks, TER remains completely stable (15.59% for FEMNIST, 10.35% for CelebA). Even at compression ratios exceeding 13,000:1 ($k=500$ for FEMNIST’s 6.6M parameters), robustness is fully preserved. This insensitivity stems from Count Sketch’s distance preservation guarantees: sketch-based distances remain sufficiently accurate to distinguish Byzantine updates from honest behavior across a wide range of compression levels.
5.5.2 Topology Robustness
Table 2 presents robustness across five network topologies for the three similarity-based methods (non-robust D-FedAvg and KRUM consistently fail at $\sim$63% TER as shown in Figure 2). SketchGuard matches BALANCE across all topologies, confirming that sketch compression does not interact negatively with network structure. In well-connected topologies (ER $p\geq 0.45$, Fully Connected), all methods maintain TER below 20% even at 80% Byzantine. In sparse Ring topology, all methods exhibit graceful degradation at 60%+ attack intensity due to limited honest neighbors—a fundamental limitation of local aggregation rather than any method-specific weakness.
Table 2: Test Error Rate (%) across topologies for robust methods, averaged over datasets and attacks (excl. backdoor).
Topology Byz.% UBAR BALANCE SG
Ring 20 19.0 19.9 19.9
40 18.9 19.8 19.6
60 30.2 30.6 30.3
80 64.3 63.7 64.2
ER (pp=0.2) 20 20.6 20.8 20.7
40 20.2 20.4 20.3
60 19.4 19.8 19.7
80 18.8 20.4 20.0
ER (pp=0.45) 20 20.1 20.4 20.3
40 20.4 20.1 19.9
60 19.6 19.8 19.5
80 19.2 20.3 20.2
ER (pp=0.6) 20 20.1 20.6 20.3
40 20.1 20.2 20.1
60 19.7 19.7 19.5
80 19.1 19.2 19.0
Fully Conn. 20 20.0 20.5 20.3
40 19.9 20.1 19.9
60 19.5 19.7 19.5
80 19.1 19.3 18.8
6 Conclusions
@@
We proposed SketchGuard, a framework for scaling Byzantine-robust decentralized federated learning through sketch-based neighbor screening. By decoupling filtering from aggregation, our approach reduces communication complexity from $O(d|\mathcal{N}_{i}|)$ to $O(k|\mathcal{N}_{i}|+d|\mathcal{S}_{i}|)$ while preserving the robustness guarantees of full-precision defenses.
@@
Our theoretical analysis establishes that Count Sketch compression maintains Byzantine resilience with controlled degradation bounds, introducing only a $(1+O(\epsilon))$ factor in the effective threshold. Experiments confirm that SketchGuard matches state-of-the-art robustness (mean TER deviation $\leq$0.5 percentage points) while reducing computation by up to 82% and communication by 50–70% under typical attack scenarios.
Impact Statement
This paper presents work whose goal is to advance the field of Distributed Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 4.7
@@ A.1 Setup and Notation
$\mathbf{w}_{i}^{t+1/2}=\mathbf{w}_{i}^{t}-\eta\mathbf{g}(\mathbf{w}_{i}^{t}),$ (10)
@@
A.2 Model Update Analysis
@@
$\quad+(1-\alpha)\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t}$ (11)
@@
$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2}+\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$ (12)
@@
$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})-\mathbf{w}_{i}^{t}$ (13)
@@
$\quad\times\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})$ (14)
A.3 Applying Smoothness
@@
$\quad+\frac{L}{2}\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|^{2}$ (15)
@@
$\quad\left.+\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$ (16)
A.4 Bounding the Quadratic Term
@@
$\quad+2\left\|\frac{1-\alpha}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|^{2}$ (17)
A.5 Impact of Sketch-Based Filtering
@@
$\quad\leq\gamma\exp(-\kappa t/T)\|\mathrm{CS}(\mathbf{w}_{i}^{t+1/2})\|$ (18)
@@
$\quad\leq\gamma\sqrt{\frac{1+\epsilon}{1-\epsilon}}\exp(-\kappa t/T)\|\mathbf{w}_{i}^{t+1/2}\|$ (19)
@@
$\left\|\frac{1}{|\mathcal{S}_{i}|}\sum_{j\in\mathcal{S}_{i}}(\mathbf{w}_{j}^{t+1/2}-\mathbf{w}_{i}^{t+1/2})\right\|\leq\gamma_{\text{eff}}\psi$ (20)
@@ A.6 Taking Expectation
$\quad+L(1-\alpha)^{2}\gamma_{\text{eff}}^{2}\psi^{2}$ (21)
A.7 Simplifying with Parameter Constraints
@@
$\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (22)
@@
A.8 Applying Strong Convexity
@@
$\|\nabla F(\mathbf{w}_{i}^{t})\|^{2}\geq 2\mu(F(\mathbf{w}_{i}^{t})-F(\mathbf{w}^{*}))$ (23)
@@
$\quad+L\eta^{2}\delta^{2}+2\gamma_{\text{eff}}\psi\rho(1-\alpha)$ (24)
A.9 Telescoping and Final Bound
@@
$\quad+\frac{2L\eta\delta^{2}}{\mu}+\frac{2\gamma_{\text{eff}}\rho\psi(1-\alpha)}{\mu\eta}$ (25)
This completes the proof of Theorem 4.7. The key observation is that the compression error only affects the convergence through $\gamma_{\text{eff}}=\gamma\sqrt{(1+\epsilon)/(1-\epsilon)}$, introducing a controllable degradation factor while maintaining the same convergence rate as the state-of-the-art.
Appendix B Proof of Theorem 4.8
@@
B.1 Starting from Smoothness